Circular No. 6327

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

V348 PUPPIS R. Baptista, University of St. Andrews; J. Patterson, Columbia Uni-

versity; D. O’Donoghue, University of Cape Town; D. Buckley, South African Astronomical Observatory; F. Jablonski, Laboratório Nacional de Astrofísica, Itajubá; T. Augusteijn, European Southern Observatory; and W. Dillon, Missouri City, TX, communicate: “We discovered that the ephemeris of Tuohy et al. (1990, Ap.J. 359, 204) for mid-eclipse times of the cataclysmic variable V348 Pup (1H 0709–36) has accumulated an error that now amounts to −0.33 cycle (−49 min). In order to revise the ephemeris, we collected a total of 44 eclipse observations, covering the period 1990 Mar. 1–1996 Jan. 27. The complete list of timings will be published in a forthcoming paper. These timings, together with the six of Tuohy et al., yield the linear heliocentric ephemeris (and 1σ uncertainties): T_{mid} = HJD 2447210.12110(14) + 0.101838911(12)E, with a dispersion of 0.0043 cycle (38 s) and a reduced χ^{2} = 1.25.”

NOTICE TO SUBSCRIBERS Effective immediately, payment for these Circulars can be accepted by

credit card (MasterCard, Visa, and Discover), subject to a $20 minimum charge. New accounts (and renewals of expired accounts) will not become effective until the first of the next month, regardless of payment date. Such orders may be placed by e-mail (to iausubs@cfa.harvard.edu), by fax (to USA 617-495-7231), or by telephone (to USA 617-495-7280, generally between 14^{h} and 21^{h} UT, Monday–Friday). When sending orders by fax or e-mail please include the following information: name (as given on the credit card); card type; credit-card number and expiration date; address at which the card is registered. If the payment is for the renewal of a current or expired account, please include your 5-digit account number. Payment for line charges and for the Catalogue of Cometary Orbits can also be made by credit card.

COMET 116P/WILD 4 Total visual magnitude estimates: Feb. 8.14 UT, 13.1 (A. Hale, Cloud-

croft, NM, 0.41-m reflector); 18.00, 12.4 (J. M. San Juan, Teruel, Spain, 0.44-m reflector); 24.78, 12.2 (K. Hornoch, Lelekovice, Czech Republic, 0.35-m reflector).

1996 February 27  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green